Suppose an ANOVA has three treatment conditions, and n = 12 in each condition, which pair of values correctly identify the df for the F-ratio?

Group of answer choices

3, 36

2, 33

2, 12

2, 36
Answer #1

The df for the F-ratio is = (k - 1, n - k). Here, k = no. of treatment conditions = 3, n = sample size = 12 * 3 = 36.
Hence, the df = (2, 33).